OpenStudy (anonymous):

HELP!!! Find the Least Common Multiple of X^2 - 5x + 6 and X^2 - 8x +15

OpenStudy (anonymous):

try factoring it

mathslover (mathslover):

Ok wait lemme explain u

OpenStudy (anonymous):

so its X -5 X-3 and x -2 ?

mathslover (mathslover):

Find the Least Common Multiple of X^2 - 5x + 6 and X^2 - 8x +15 \[\large{x^2-5x+6}\] \[\large{x^2-3x-2x+6}\] \[\large{x(x-3)-2(x-3)}\]

mathslover (mathslover):

that is : \(\large{(x-2)(x-3)}\)= \(\large{x^2-5x+6}\)

mathslover (mathslover):

similarly: \(\large{x^2-8x+15=x^2-5x-3x+15}\) \(\large{=x(x-5)-3(x-5)=(x-3)(x-5)}\)

mathslover (mathslover):

hence we have least common multiple : (x-3)

mathslover (mathslover):

got it @Lunarwolf58 ?